cdiate the ter-tax cost of overlapping nteest tax bracket. ompany, an American company, is contemplating offering a new $50 million bond is standing $50 million bond issue. The company wishes to take advantage of the to replace an out decline in interest rates that has occurred since the initial bond issuance. are described in what follows. The company is in the 40% tax bracket. old bonds. The outstanding bonds have a $1,000 face value and a 9% coupon interest rate. hey wer of $1,000, and the compan New bonds. The new bonds would have a $1,000 face value, a 7% coupon interest rate, and a 15-year maturity. They could be sold at their face value. The flotation cost of the new bonds would be $500,000. The company does not expect to have any overlapping interest. a Calculate the tax savings that are expected from the unamortised portion of the old bonds flotation cost. e issued five years ago with a 20-year maturity. They were initially sold for their face value y incurred $350,000 in flotation costs. They are callable at $1,090 e annual tax savings from the flotation cost of the new bonds, assuming the 15-year amortisation.
14: Long-Term Debt and Leasing c Calculate the after-tax cost of the call premium that is required to retire the old bonds d Determine the initial investment that is required to call the old bonds and issue the new e Calculate the annual cash flow savings, if any, which are expected from the proposed refunding decision bonds. bond f If the company has a 42% after-tax cost of debt, find the net present value (NPV) of the bond refunding decision. Would you recommend the proposed refunding? Explain your answer. P14-5 Web Tools Company is considering using the proceeds from a new $50 million bond issue to ca and retire its outstanding $50 million bond issue. The details of both bond issues are outlined in what follows. The company is in the 40% tax bracket. old bonds. The company's old issue has a coupon interest rate of 10%, was issued four years ago, and had a 20-year maturity. The bonds sold at a $10 discount from their $1,000 face value, flotation costs were $420,000, and their call price is $1,100. New bonds. The new bonds are expected to sell at par ($1,000), have a 16-year maturity, and have flotation costs of $520,000. The company will have a two-month period of overlapping interest while it retires the old bonds. What is the initial investment that is required to call the old bonds and issue the new bonds? b What are the annual cash flow savings, if any, from the proposed bond refunding decision if the new bonds have an 8% coupon interest rate? If the new bonds have a 9% coupon interest rate? a c Construct a table showing the net present value (NP) of refunding under the two circumstances given in part (b) when (1) the company's after-tax cost of debt is 4.8% 1008 3 (1-0401 and(2 this cost is 5.4% [0.09 3 (1-040). d Given the circumstances described in part (c), discuss when refunding would be favourable and when it would not e If the two circumstances summarised in your answer to part (d) were equally probable (each had bability of 25%), would you recommend refunding? Explain your answer. ntermine the yearly after-tax
Answer #1
a Tax savings expected from unamortised portion of Flotation cost of old Bond
Flotation cost incurred $350,000
Number of years to maturity of old Bond 20
Annual amortization(350000/20) $17,500
Unamortised flotation cost =17500*15= $262,500
Tax savings on unamortised flotation cost = $105,000 (262500*40%)
b. Flotation cost of the new issue $500,000
Number of years to maturity 15
Annual amortization=500000/15= $33,333.33
Tax rate 40%
Annual tax saving on amortization of new bond floatation cost $13,333 (33333.33*40%)
c After tax cost of the Call Premium for retiring the old bonds
Number of Bonds outstanding=50million/1000                 50,000
Call premium per bond(1090-1000) $90
Before tax call premium=50000*90= $4,500,000
After tax call premium =4500000*(1-40%) $2,700,000
d Initial Investments Required:
After tax Call premium -$2,700,000
Flotation cost of new bond ($500,000)
Tax savings on unamortised flotation cost = $105,000
Net Initial Investment required -$3,095,000 (Cash outflow , hence with negative sign)
e Annual Cash Flow savings;
Savings on interest:
Savings on Annual Coupon on old bond=50 million*9%= $4,500,000
Payment of Annual Coupon on new bond=50 million*7%= -$3,500,000
Net before tax interest savings $1,000,000
After tax cash flow on savings of interest $600,000 (1000000*(1-40%)
Savings on tax shield on flotation cost amortization
Tax shield on amortization of flotation cost of new bond $13,333
Tax shield lost on amortization of flotation cost of old bond ($7,000) (17500*40%)
Net Flotation cost tax savings $6,333
Annual Cash Flow savings for 15 years=600000+6333= $606,333
f NET PRESENT VALUE
I Initial investment -$3,095,000
Rate Annual after tax cost of debt 4.20%
Nper Number of years of cash flow 15
Pmt Annual Cash Flow savings $606,333
PV Present value of annual cash flow savings $6,648,138 (Using PV function of excel with Rate=4.2%, Nper=15, Pmt=-606333)
Excel Command : PV(4.2%,15,-606333)
NPV=PV+I NET PRESENT VALUE=6648138-3095000= $3,553,138
Yes , We would recommend the proposal of refunding
NPV is Positive